Database last updated on Mon, 02 Feb 2026 03:40:49 +0000 by localhost
Instrument constant changes for the week before 2023238 (3 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 29 rows)
Measurements and tests summary.
_jday
_loc
_serial
_skc
aode
ds
dt
ed
fz
hg
rs
sl
uv
zs
238
GOOSE BAY
#018
ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.14701)
4
4
1
2
4
18
1
6
63
85
(0 sec)
Dead time (0 sec)
File name
Count
Serial
Expected (+/-5)
Measured
30-day history
B23823.018
1
#018
35
35